OpenStudy (anonymous):

4c^2=7c The solution set is?

OpenStudy (anonymous):

\[4c^2=7c\]\[4c^2-7c=0\]\[c(4c-7)=0\] 1. if c = 0 , 4c² − 7c = 0 2. if 4c−7 = 0, 4c² − 7c = 0 \[4c−7 = 0\]\[4c=7\]\[c=\frac{7}{4}\] Answer: \[c=0 , c=\frac{7}{4}\]

OpenStudy (anonymous):

oh ok..because i ended up with 0, 1.75

OpenStudy (anonymous):